OpenStudy (anonymous):

what is the solution of the matrix equation { 5, -2 2,1]x= [4 -4] A) [2 -4] B) [-10 -24] C) [-10 24] D) [10 24]

hero (hero):

I assume this is your matrix equation

OpenStudy (e.mccormick):

Or perhaps: \(\begin{bmatrix} 5 & -2\\ 2 & 1 \end{bmatrix}x=\begin{bmatrix} 4\\ -4 \end{bmatrix}\)

hero (hero):

It is in the form \(Ax = B\) Find the inverse of matrix A then take the inverse of both sides to isolate x: \(x = A^{-1}B\)

hero (hero):

Then multiply the inverse of A with B to get your final result.

OpenStudy (anonymous):

wow sorry i messed up that 1 is a -1 and the positive 4 is a 2 and yes that is my matrix

hero (hero):

\[\begin{bmatrix} 5 & -2\\ 2 & -1 \end{bmatrix}x=\begin{bmatrix} 2\\ -4 \end{bmatrix}\]

OpenStudy (anonymous):

yes perfect now how do you the next step sorry I'm not very good with matrixes

hero (hero):

\[A^{-1} = \frac{1}{|A|} \begin{bmatrix} d & -b \\ -c & a\end{bmatrix}\]

OpenStudy (anonymous):

so would the answer be c [-10] [ 24]

OpenStudy (e.mccormick):

What Hero is using is a principal of matrices that is based off of the multiplicative inverse. \(5\cdot \dfrac{1}{5}=1\) That is true for pretty much any number and also works in matrices. So what he is saying is that your original form is: \(Ax=b\) You need to cancel the \(A\) on the left, but that will make a change in the right. To do so, you need the multiplicative inverse: \(Ax=b\implies \) \(A^{-1}Ax=A^{-1}b\implies \) \(1x=A^{-1}b\implies \) \(x=A^{-1}b\) So the first step in this is to find \(A^{-1}\).

OpenStudy (anonymous):

you guys are like genius compared to me in this so is a -4?

OpenStudy (e.mccormick):

We have just done this before. Then he showed you a form of this: If \(A=\begin{bmatrix} a & b\\ c & d \end{bmatrix}\) then: \(A^{-1}=\dfrac{1}{ad-bc}\begin{bmatrix} d & -b\\ -c & a \end{bmatrix}\) That help?

OpenStudy (anonymous):

damn I'm sorry so its 1/ (-1)(5) - (-2)(2)? @e.mccormick

OpenStudy (e.mccormick):

For that part, yes. Then you multiply through the modified matrix (adjoint matrix) with that.

OpenStudy (anonymous):

man i feel so dumb could you show me how to do it

OpenStudy (e.mccormick):

Well, what did \(\dfrac{1}{(-1)(5) - (-2)(2)}\) get you?

OpenStudy (anonymous):

okay so you -5 - (-4)= 9

OpenStudy (e.mccormick):

Ummm... When you distribute the \(-(-4)\), what do you get?

OpenStudy (anonymous):

-1 sorry

OpenStudy (e.mccormick):

OK. Good! So we are at: \(A^{-1}=-1\begin{bmatrix} d & -b\\ -c & a \end{bmatrix}\) So now, the a, b, c, and d parts. You take these from the original matrix: \(A=\begin{bmatrix} a & b\\ c & d \end{bmatrix}\) So, what are they?

OpenStudy (anonymous):

[-5 2] [ -2 1 ] ?

OpenStudy (e.mccormick):

OK, you did the sign swap, but look at the position of a and d. If you learned about a cofactor matrix, this is all about that.

OpenStudy (anonymous):

okay so (-1 2) (-2 5)

OpenStudy (e.mccormick):

Just a little FYI in case my terms are confusing you: Depending on where you are with linear algebra, or just algebra, they may throw all these terms at you or not. If you are learning this as part of algebra, and not a linear class, al they usually talk about are: matrix transpose matrix cofactor matrix inverse matrix determinant Then they just give you a few set formulas to remember. If you are taking linear algebra, they then say: adjoint/adjunct/adjugate matrix = transpose of cofactor matrix eigen values eigen vectors column vectors row vectors pivots/free variables So if I do not know where you are, I might use some terms you are unfamiliar with.

OpenStudy (anonymous):

thats okay I've seen these before matrix cofactor matrix inverse matrix

OpenStudy (e.mccormick):

OK. So, here is where you are at: \(A^{-1}=-1\begin{bmatrix} d & -b\\ -c & a \end{bmatrix}=-1\begin{bmatrix} -1 & 2\\ -2 & 5 \end{bmatrix} \)

OpenStudy (e.mccormick):

So, multiply through the -1 (swap the signs).

OpenStudy (anonymous):

[1 - 2] [2 -5] right

OpenStudy (e.mccormick):

Yes. So, do you understand what you need to do with that now? How to plug it in?

OpenStudy (anonymous):

no really, I'm really sorry man i know its probably a pain for you to do this

OpenStudy (anonymous):

not*

OpenStudy (e.mccormick):

It is just a scroll up at this point. OK, we are at: \(A^{-1}=\begin{bmatrix} 1 & -2\\ 2 & -5 \end{bmatrix} \) Now, remember, this was just to find what is needed for: \(x=A^{-1}b\) From the original, \(b=\begin{bmatrix}2\\ -4 \end{bmatrix}\) and you just found \(A^{-1}\) so: \(x=\begin{bmatrix}1 & -2\\ 2 & -5\end{bmatrix}\begin{bmatrix}2\\ -4 \end{bmatrix}\) Can you do that now? Oh, and how I Hero and I did the whole formatted things is with \(\LaTeX\) `\(x=\begin{bmatrix}1 & -2\\ 2 & -5\end{bmatrix}\begin{bmatrix}2\\ -4 \end{bmatrix}\)`

OpenStudy (anonymous):

okay i see how you did the format thing and would it be the answer D

OpenStudy (e.mccormick):

Yep. That would be it in the end. Here is a reference to that whole inverse thing we used. It should be similar to what you see in an algebra book: http://www.mathwords.com/i/inverse_of_a_matrix.htm If you can get this process down, and it really is just a process, you should do fine on any test.

OpenStudy (anonymous):

yesss! Thanks a lot man i really appreciate it, your a legend

OpenStudy (e.mccormick):

Hero is a legend. I just helped out a little.

OpenStudy (anonymous):

haha both you guys are, but anyways i really appreciate it, thanks

OpenStudy (e.mccormick):

np. Have fun!

OpenStudy (anonymous):

hah il try too